The following information is known for a given system: 1. The total System Operating time for a designated period is 10,000 hours 2. The corrective maintenance rate is .004 maintenance action per hour of system operation 3. The perventive maintenance rate is .001 maintenance action per hour of system operation 4. The average downtime per maintenance action is 50 hours with a mean downtime of 5 hours for corrective maintenance

a. Calculate the following: MTBF, MTBM, Ai, Ao?

b. What is the relability for 100 hours of operation

c. Assume that the operational availability requirement for the system is .8. Lets say that the actual demonstrated MTBM is 100 hour which is less than the calculated value above. Given this condition, what is the new Ao value? What corrective action action needs to be taken to maintain mission availability requirements (in other words retain an Ao value of .80)?
Answer #1

Answer A;- In the question the data is given as below:-

Operating time=10,000 hrs.

Corrective maintenance rate =0.004

Therefore total correct maintenance = 10000*0.004=40

ʎ = total number of failures / total operating hours

ʎ = 40 / 10000 =0.004

rate of preventive maintenance =0.001

So total corrective maintenance =10000*0.001 =10

Total number of maintenance = 40+10 =50

MTBF= 1/ ʎ = 1/0.004 =250 hours

Total down time during corrective maintenance =total corrective maintenance * average downtime =40*5=200

Average Downtime for preventive maintenance=50-5=45 hours

total down time for preventive maintenance = 45*10 = 450 hours

Total downtime = 450+200 = 650 hours

Up time = 10000-650 = 9350 hours

MTBM = uptime / total maintenance   =9350/50=187

MTTR= total down time / total maintenance   =650/40=16.25 hr

Ai= MTBDF / (MTBF+MTTR) = 250/(250+16.25) =0.939

Ao = MTBM/(MTBM+MDT) =187/(187+50) =0.789

Answer b;- For 100 hour operation, reliability= e-λt=e(-0.04*100)=0.67

Answer C: In the question

MTBM= 100

MDT= 50

Ao = MTBM/(MTBM+MDT) =100/(100+50) =0.667

Thus it is clear that we have to decrease MDT to improve the value of Ao.